OpenStudy (anonymous):

What is the slope of the line y = -2/3x + 3

OpenStudy (anonymous):

i think its 3

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

m = 3 m = -2/3 m = 2/3 m = -3

OpenStudy (anonymous):

awesome. i would have gotten it wrong

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

no problem